OpenStudy (anonymous):

converting 29 ounces to pounds? needed for apex.

OpenStudy (espex):

You should know that \(\frac{1lb}{16oz}\) Writing it like this will allow you to do a differential analysis and when your terms cancel to leave you the desired units, all that is left is the math.

OpenStudy (anonymous):

ole ufa

OpenStudy (anonymous):

im so confused can you put it in english please @eSpeX

OpenStudy (espex):

If you are converting ounces to pounds, and you're doing a differential analysis, you are trying to get your units to cancel out--just as you would reducing any fraction. SO, you have:\(\frac{29~oz}{1}\) that you wish to convert to \(\frac{pounds}{1}\)

OpenStudy (anonymous):

oh i get it now thank you @eSpeX

OpenStudy (espex):

You're welcome.